OpenStudy (anonymous):

What are the atomic number and the atomic mass of the element formed when 84 218Po emits a beta particle? What are they if the polonium emits an alpha particle? @Abhisar

OpenStudy (anonymous):

\[\frac{ 218 }{84 }Po\]

OpenStudy (abmon98):

In Beta decay a neutron changes into a proton plus an electron. The proton stays in the nucleus and the electron leaves the atom with high energy, and we call it a beta particle. When a beta particle is emitted from the nucleus the nucleus has one more proton and one less neutron. This means the atomic mass number remains unchanged and the atomic number increases by 1. Atomic Numer=84+1=85

OpenStudy (abmon98):

Alpha decay: the nucleus looses two protons and neutrons, so the atomic mass number will decrease by 4 and the atomic number decreases by 2. atomic mass number: 218-4=214 atomic number: 84-2=82

OpenStudy (anonymous):

@Abhisar